The following function is continuous at x = 0:


𝑓(π‘₯) =


(1βˆ’cosπ‘˜π‘₯)/(xtanx)


, π‘“π‘œπ‘Ÿ π‘₯ β‰  0 π‘Žπ‘›π‘‘ 𝑓(0) = 3. find k.

1
Expert's answer
2022-04-17T10:08:40-0400

Using the well-known trigonometric identity, we get: "1-\cos\,kx=2\,sin^2\frac{k}{2}x". Thus, the function can be rewritten as: "\frac{2\sin^2\frac{k}{2}x}{x\,\tan\,x}". Remind the well-known limit: "\frac{\sin\,x}{x}\rightarrow1,x\rightarrow0". We rewrite the expression and receive: "\frac{xk^2}{4}\frac{2\sin^2\frac{k}{2}x}{\frac{k^2}{4}x^2\,\tan\,x}". Since, "\frac{x}{\tan\,x}\rightarrow1,x\rightarrow0" and "\frac{\sin\,x}{x}\rightarrow1,x\rightarrow0", we get: "\frac{xk^2}{4}\frac{2\sin^2\frac{k}{2}x}{\frac{k^2}{4}x^2\,\tan\,x}\rightarrow\frac{k^2}{4},x\rightarrow0". From the latter and from the formulation of the task we get: "\frac{k^2}{4}=3". Finally, we get: "k=\pm\sqrt{12}." Thus, the answer is: "k=\pm\sqrt{12}."